OpenStudy (anonymous):

Find real or imaginary solutions of the equation by factoring x^4-10x^2-9=0 Find real or imaginary solutions of the equation by factoring x^4-10x^2-9=0 @Mathematics

OpenStudy (anonymous):

Let x^2 be y x^4 =y^2 replacing them in equation we get y^2-10y-9=0

OpenStudy (anonymous):

bleep, are you sure this is a -9?

OpenStudy (anonymous):

i almost did the same thing, sumanthy!!!

OpenStudy (anonymous):

Yes , I just deleted the post as I checked it -9 :) Thanks Mandolino

OpenStudy (anonymous):

bleep are you checking?

OpenStudy (anonymous):

my fail positive 9

OpenStudy (anonymous):

\[x^4-10x^2-9=0\]factor as a trinomial: either use the substitution that sumanthy suggested or just write x^2 where you would usually write x\[(x^2-9)(x^2-1)=0\]factor each binomial as a difference of two squares\[(x-3)(x+3)(x-1)(x+1)=0\]set each factor equal to zero to solve\[x-3=0 \cup x+3=0 \cup x-1=0 \cup x+1=0\]solution set is\[\left\{ -3,-1,1,3 \right\}\]

OpenStudy (anonymous):

When you have trinomial where the ratio of the exponents is 4:2 (as we have here) or 2:1, then you can use quadratic equation methods. So I just looked for factors of +9 that sum to -10; those are -9 and -1; this is FOIL in reverse

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

yeah i get it now thanks. we'll see if i understand the harder ones but i understand the basics

OpenStudy (anonymous):

cool! :})